# fMRI-3D: A Comprehensive Dataset for Enhancing fMRI-based 3D Reconstruction

### v1

@@ 8 Conclusion
In this paper, we introduce the innovative task of Recon3DMind and its accompanying large-scale dataset, fMRI-3D, across various settings for the first time. Technologically, we develop a novel three-stage framework that involves multiple brain regions, including those associated with human 3D vision, specifically designed for this task. This approach establishes new benchmarks in the field and demonstrates the feasibility of the task. Initially, our model proficiently extracts features from fMRI frames. In the second stage, it employs a diffusion module to transition these fMRI features into the vision domain. Finally, in the third stage, the vision features are transformed into 3D models using an advanced 3D generation model. Our comprehensive experimental results and analyses affirm the model’s effectiveness in accurately extracting fMRI features and converting them into their corresponding 3D objects. We also conduct an in-depth analysis of our proposed fMRI-3D dataset and the features extracted by MinD-3D, further validating both the quality of fMRI-3D and the effectiveness of MinD-3D. This pioneering work not only opens up a new avenue in neuroimaging and 3D reconstruction but also paves the way for future research aimed at a deeper understanding and visualization of neural representations in 3D vision.

### v3

@@ 8 Conclusion
In this paper, we introduce the innovative task of Recon3DMind with texture, alongside the first large-scale dataset—fMRI-3D—across various settings. Technologically, we present a novel end-to-end framework, MinD-3D++, which integrates multiple brain regions, including those associated with human 3D vision, specifically designed for this task. This approach not only establishes new benchmarks in the field but also demonstrates the feasibility of reconstructing textured 3D objects from human brain data. Our model begins by proficiently extracting features from fMRI signals using a contrastive learning loss. It then generates multi-view images of target objects, which are subsequently used to reconstruct textured 3D models. Comprehensive experimental results and analyses confirm the effectiveness of MinD-3D++ in accurately extracting fMRI features and converting them into their corresponding 3D objects. Additionally, we perform an in-depth analysis of our proposed fMRI-3D dataset and examine the features extracted by MinD-3D. These evaluations further validate both the quality of the fMRI-3D dataset and the effectiveness of our approach. This pioneering work not only opens a new avenue in neuroimaging and 3D reconstruction but also paves the way for future research aimed at a deeper understanding and visualization of neural representations in 3D vision.